# DeltaEdit: Enhancing Sequential Editing in Large Language Models by Controlling Superimposed Noise

### v1

@@ Abstract
Sequential knowledge editing techniques aim to continuously update the knowledge in large language models at a low cost, preventing the models from generating outdated or incorrect information. However, existing sequential editing methods suffer from a significant decline in editing success rates after long-term editing. Through theoretical analysis and experiments, we identify that as the number of edits increases, the model’s output increasingly deviates from the desired target, leading to a drop in editing success rates. We refer to this issue as the accumulation of superimposed noise problem. To address this, we identify the factors contributing to this deviation and propose DeltaEdit, a novel method that optimizes update parameters through a dynamic orthogonal constraints strategy, effectively reducing interference between edits to mitigate deviation. Experimental results demonstrate that DeltaEdit significantly outperforms existing methods in edit success rates and the retention of generalization capabilities, ensuring stable and reliable model performance even under extensive sequential editing.
@@ 1 Introduction
Large language Models encode vast amounts of knowledge acquired during pretraining[1; 2; 3]. This makes them invaluable as knowledge bases across various applications[4; 5]. However, they are prone to generating inaccurate or outdated information[6; 7], necessitating continuous updates to maintain their accuracy and reliability. While fine-tuning offers a potential solution for knowledge updates, it is computationally expensive and risks catastrophic forgetting[8; 9]. To overcome these limitations, knowledge editing techniques have emerged as an efficient alternative. They enable precise updates with minimal computational cost, while preserving the integrity of existing knowledge[10; 11; 12]. These techniques can be broadly divided into two categories: parameter-preserving methods[9; 13; 14; 15; 16; 17], which incorporate external modules without altering the model’s parameters, and parameter-modifying methods[18; 19; 10; 20; 21], which directly adjust the model’s parameters. This study focuses on the latter, parameter-modifying methods.
Current mainstream editing methods adopt the locate-then-edit paradigm[10; 22; 23], which involves first identifying the most impactful model parameters $W$ and then introducing update parameters $\Delta$ to perform the desired edit. While these methods excel in single-edit tasks, real-world applications often demand multiple consecutive updates to accommodate rapidly evolving knowledge. This shift highlights the importance of sequential editing, which demands performing a series of edits while ensuring that all updated knowledge is accurately integrated and retained. Prior studies[24] have shown that naively extending single-edit methods to sequential tasks can lead to reduced edit success rates and degradation in model performance. To address these issues, researchers have explored several critical factors[22; 23; 21]. These efforts have led to the development of new methods specifically designed for sequential editing.
Despite recent advances, most existing studies treat the update parameters for editing as a monolithic entity, neglecting the fine-grained dynamics of the update process and the potential interactions between successive edits. In this paper, we conduct a comprehensive investigation into the dynamic behavior of sequential editing. We identify a critical phenomenon: as the number of edits increases, the model’s output increasingly deviates from its intended target. We term this cumulative deviation superimposed noise. Through experiments, we demonstrate the adverse effects of superimposed noise on model performance.
@@
In this paper, we propose DeltaEdit, a novel sequential editing method designed to mitigate the effects of superimposed noise. DeltaEdit introduces a dynamic orthogonal constraint strategy to explicitly optimize influence vectors during the editing process, reducing interference between updates. Experimental results demonstrate that DeltaEdit effectively reduces superimposed noise, enhances the precision of edits, and maintains the model’s generalization capability.
To summarize, the main contributions of this work are as follows:
@@
    We uncover and define superimposed noise as a core limitation in sequential editing tasks. Furthermore, through experiments on multiple models and editing methods, we demonstrate its negative impact on the performance of the edited models.
@@
    We further analyze the factors affecting superimposed noise. By understanding these factors, we identify strategies to reduce superimposed noise and improve the performance of sequential editing methods.
@@
    We develop DeltaEdit, an innovative sequential editing method that incorporates a dynamic orthogonal constraint strategy. Through extensive experiments, we demonstrate that DeltaEdit significantly outperforms existing methods in terms of edit success, retention of general capabilities, and overall robustness.
@@ 2.1 Knowledge Editing
From the perspective of whether to modify model parameters, [11] categorizes knowledge editing methods into two major types: parameter-preserving methods and parameter-modifying methods. This paper primarily focuses on the latter. One line of work employs meta-learning approaches to edit language models through a hypernetwork. KE[18] utilizes a bidirectional LSTM to predict weight updates for editing, whereas MEND[7] applies a low-rank decomposition of gradients to fine-tune the language model. Another line of work, based on conclusions drawn from causal probes [25; 10], performs edits in the feed-forward networks of middle model layers. KN[25] achieves knowledge editing by modifying the activation values of specific neurons. ROME[10] uses normal equations to compute the update parameters required for editing, and MEMIT[20] further extends this approach to support batch editing.
@@ 2.2 Sequential Editing
Gupta et al.[24] points out that performing multiple consecutive edits can lead to model performance degradation. Yang et al.[26] identified perplexity as an effective metric for detecting model collapse during sequence editing. Gupta et al.[27] conducted an analysis demonstrating that the utilization of two distinct types of key vectors in ROME contributes to model collapse in sequence editing. Yang et al.[28] reached similar conclusions and further clarified that the distributional discrepancies in key vectors, caused by the use of different types of key vectors, are a key factor driving model collapse. By analyzing update parameters, Hu et al.[29] finds that the overlap of input representations in the whitening space is a key factor contributing to poor editing performance. Ma et al.[22] theoretically analyzes that the bottleneck limiting sequential editing in models lies in the condition number of matrices and proposes PRUNE, which supports sequential editing by controlling the growth of the condition number. Gu et al.[23] observes that editing causes excessive parameter changes and proposes RECT, which improves editing performance by sparsifying update parameters. Fang et al.[21] proposes AlphaEdit, which achieves nearly lossless sequential editing by restricting the solution space of update parameters to a specific null space.
3 Analysis on Bottleneck of Sequential Editing
@@ 3.1 Preliminary
Autoregressive Language Model  Autoregressive language models[30] generate the next token based on the preceding tokens. Modern autoregressive models are predominantly composed of multiple stacked transformer layers. In such models, the hidden state of the $l$ layer, denoted as $h^{l}$, is calculated as:
@@
Here, $a_{l}$ is the output of the attention block, $m_{l}$ is the output of the feed-forward network (FFN), $\sigma$ is the activation function, and $\gamma$ denotes layer normalization. Studies[31] show that the parameter matrices of the FFN, $W_{\text{out}}$ and $W_{\text{in}}$, encode knowledge acquired during pretraining. Building on this finding, most knowledge editing methods focus on modifying the FFN to update knowledge.
Sequential Editing  Sequential editing[32] refers to performing continuous model editing on a model. Model editing aims to modify the knowledge encoded in a model. Following previous research[10], we define knowledge in the form of triples, represented as $(s,r,o)$, where $s$ is the subject, $r$ is the relation, and $o$ is the object. For example, if a model memorizes the knowledge triple $(s=\text{iPhone},r=\text{Latest model},o=\text{iPhone 16})$, providing a language prompt $p(s,r)$, such as “The latest model of iPhone is,” as input causes the model to output “iPhone 16.” Model editing refers to the process of replacing the original knowledge triple $(s,r,o)$ with a new knowledge triple $(s,r,o^{*})$, and we denote this operation as $\mathcal{E}=(s,r,o,o^{*})$. In a sequential editing task of length $T$, where an editing sequence $\mathbb{E}_{T}=(\mathcal{E}_{1},\mathcal{E}_{2},\dots,\mathcal{E}_{T})$ is performed, each edit $\mathcal{E}_{i}$ builds upon the results of the previous edit $\mathcal{E}_{i-1}$.
Current knowledge editing methods[10; 20; 21] primarily adopt the locate-then-edit paradigm to implement model editing. For a model $M$, these methods first locate the most appropriate parameters $W$ within the model, and then update them using the following approach:
@@ 3.2 Update Parameter
Most knowledge editing methods solve for update parameter $\Delta$ by applying the normal equation. For example, MEMIT[20] leverages the knowledge storage properties of FFN layers, identifies the $W_{out}^{l}$ of a suitable layer $l$ as the target for editing, and proposes the following equation:
@@
AlphaEdit[21] has a similar solution:
@@
$\Delta=\alpha\beta^{\top},\alpha=R$ (9)
@@
Figure 1: The changes in Efficacytop (the rate of successfully maximizing the object’s probability) and $noise_{E}$ with the number of edits. The top figure displays results for GPT2-XL using AlphaEdit and MEMIT, while the bottom figure shows results for LLaMA3-8B using only AlphaEdit. MEMIT is excluded for LLaMA3-8B due to its excessively high $noise_{E}$, making visualization difficult.
@@ 3.3 The Problem of Superimposed Noise
Superimposed Noise  In the sequential editing task, a series of editing operations is performed on the model. Upon completing the entire sequence of edits $\mathbb{E}_{T}$, the parameters $W_{out}^{l}$ of model’s $l$-th layer are updated. For the representation $k_{e}$ of the subject token being edited during operation $\mathcal{E}_{e}$, the output deviates from the original $W_{out}^{l}k_{e}$. This deviation can be formally analyzed using the $L_{2}$ norm:
@@
Here, $\|W_{out}^{l}k_{e}\|_{2}$ is a constant, while $\|\sum_{i}\Delta_{i}k_{e}\|_{2}$ determines the upper bound of the deviation. In editing operation $\mathcal{E}_{e}$, the update parameter can be expressed as $\Delta_{e}$. Ideally, after multiple sequential edits, different editing operations should remain independent and not interfere with one another. Under such ideal conditions, the deviation associated with $\mathcal{E}_{e}$ satisfies:
@@
However, in practice, interference between different editing operations leads to a deviation. To formally quantify this interference, we refer to this deviation as noise and define the noise associated with $\mathcal{E}_{e}$ as:
@@
Here, $noise_{e}$ represents the superimposed noise experienced by $\mathcal{E}_{e}$ due to the influence of other editing operations in the sequence. To measure the overall superimposed noise level of all editing operations, we further define the average $noise_{E}$ of all editing operations as:
@@
The Impact of Noise on Editing Performance  To investigate the impact of superimposed noise on editing performance, experiments are conducted using AlphaEdit and MEMIT on GPT2-XL[33] and Llama3-8B[34]. In these experiments, the models are sequentially edited using the CounterFact dataset[10]. Details of the experimental setup are provided in the appendix A. The results presented in Figure 1 reveal the following trends:
@@
These findings underscore the importance of controlling superimposed noise in sequential editing tasks. Unchecked noise accumulation undermines the effectiveness of editing operations and can significantly impair the edited model. Therefore, mitigating superimposed noise is critical for ensuring robust and reliable model behavior.
4 DeltaEdit
As analyzed in Section 3.3, the progressive increase in superimposed noise during sequential editing degrades the performance of the edited model. To address this issue, we conduct a more in-depth analysis of the factors underlying noise. Building on these insights, we propose a novel editing approach, DeltaEdit, to effectively reduce superimposed noise and improve the reliability of the edited model.
Figure 2: The variation curve of $k^{\top}\beta$ for AlphaEdit and MEMIT as the number of edits increases, with GPT2-XL serving as the edited model. As the results of Llama3-8B are not suitable for display in a line chart, they are presented in Appendix C.1.
@@ 4.1 Factors Affecting Superimposed Noise
The experimental results are shown in Figure 2. The results indicate that the $k^{\top}\beta$ values obtained using AlphaEdit are significantly smaller than those of MEMIT. This improvement is primarily attributed to AlphaEdit’s use of null space projection and the inclusion of $K_{p}$ in the computation of $\beta$, which reduces the overlap of information between $\beta$ and the input representations of unrelated tokens. This demonstrates that optimizing $\beta$ to minimize erroneous activations is an effective strategy. However, as the number of sequential edits increases, the performance of AlphaEdit still degrades significantly. This suggests that simply reducing $k_{e}^{\top}\beta_{i}$ ($i\neq e$) is not sufficient to completely resolve the noise issue in sequential editing. To further reduce superimposed noise, the other influencing factor $\alpha_{i}^{\top}\alpha_{j}$ should be given greater attention. Based on this insight, we propose a new sequential editing method, DeltaEdit.
4.2 The Implementation of DeltaEdit
To ensure that $\alpha_{i}^{\top}\alpha_{j}$ (for all $i,j,i\neq j$) approaches zero, it is essential to minimize the overlap of $\alpha$ with previous edits during training. Considering the constrained solution space for $\alpha$, we propose DeltaEdit, a method that leverages a dynamic control strategy based on historical editing information, which aims to reduce interference while ensuring editing efficiency.

### v2

# On the Superimposed Noise Accumulation Problem in Sequential Knowledge Editing of Large Language Models

@@ Abstract
Sequential knowledge editing techniques aim to continuously update knowledge in large language models at low cost, preventing models from generating outdated or incorrect information. However, existing sequential editing methods suffer from a significant decline in editing success rates after long-term editing. Through theoretical analysis and experiments, our findings reveal that as the number of edits increases, the model’s output increasingly deviates from the desired target, leading to a drop in editing success rates. We refer to this issue as the superimposed noise accumulation problem. Our further analysis demonstrates that the problem is related to the erroneous activation of irrelevant knowledge and conflicts between activated knowledge. Based on this analysis, a method named DeltaEdit is proposed that reduces conflicts between knowledge through dynamic orthogonal constraint strategies. Experiments show that DeltaEdit significantly reduces superimposed noise, achieving a 16.8% improvement in editing performance over the strongest baseline.
@@ 1 Introduction
Large language Models encode vast amounts of knowledge acquired during pretraining (brown2020language; petroni2019language; roberts2020much). This makes them invaluable as knowledge bases across various applications (thirunavukarasu2023large; alkhamissi2022review). However, they are prone to generating inaccurate or outdated information (de2021editing; mitchell2021fast), necessitating continuous updates to maintain their accuracy and reliability. While fine-tuning offers a potential solution for knowledge updates, it is computationally expensive and risks catastrophic forgetting (luo2024empiricalstudycatastrophicforgetting; mitchell2022memory; huang2025selfaug). To overcome these limitations, knowledge editing techniques have emerged as an efficient alternative. They enable precise updates with minimal computational cost, while preserving the integrity of existing knowledge (meng2023locatingeditingfactualassociations; yao2023editinglargelanguagemodels; wang2023knowledgeeditinglargelanguage; li2025mindbridge).
Current mainstream editing methods adopt the locate-then-edit paradigm (meng2023locatingeditingfactualassociations; ma2024perturbationrestrainedsequentialmodelediting; gu2024modeleditingharmsgeneral), which involves first identifying the most impactful model parameters $W$ and then introducing update parameters $\Delta$ to perform the desired edit. While these methods excel in single-edit tasks, real-world applications often demand multiple consecutive updates to accommodate rapidly evolving knowledge. This shift highlights the importance of sequential editing, which demands performing a series of edits while ensuring that all updated knowledge is accurately integrated and retained. Prior studies (gupta2024model) have shown that naively extending single-edit methods to sequential tasks can lead to reduced edit success rates and degradation in model performance. To address these issues, researchers have explored several critical factors (ma2024perturbationrestrainedsequentialmodelediting; gu2024modeleditingharmsgeneral; fang2024alphaedit). These efforts have led to the development of new methods specifically designed for sequential editing.
Despite recent advances, most existing studies on sequence editing remain superficial, they treat the update parameters for editing as a monolithic entity, neglecting the fine-grained dynamics of the update process and the potential interactions between successive edits. In this paper, we conduct a comprehensive investigation into the dynamic behavior of sequential editing. We identify a critical phenomenon: as the number of edits increases, the model’s output increasingly deviates from its intended target. As shown in Figure 1, a user query typically activates not only the correct knowledge but also numerous irrelevant knowledge. The superposition of irrelevant knowledge makes it difficult for the correct knowledge to be properly output. We refer to this issue as the superimposed noise accumulation problem. Through our experiments, we demonstrate that the accumulation of superimposed noise is a major contributor leading to decreased editing success rates and even model collapse.
@@
Based on the preceding analysis, we propose DeltaEdit, a novel sequential editing method designed to mitigate the effects of superimposed noise. DeltaEdit introduces a dynamic orthogonal constraint strategy to explicitly optimize influence vectors during the editing process, reducing interference between updates. Experimental results demonstrate that, compared to existing methods, DeltaEdit more effectively mitigates superimposed noise and achieves superior sequential editing performance. In particular, it attains a 16.8% improvement in editing performance over the strongest baseline. Furthermore, DeltaEdit better preserves model capabilities and representation distribution, thereby significantly enhancing the reliability and sustainability of sequential editing. To summarize, the main contributions of this work are as follows:
@@
    We uncover and define superimposed noise as a core limitation in sequential editing tasks. Furthermore, through experiments on multiple models and editing methods, we demonstrate that the problem of superimposed noise accumulation is a critical cause of the decline in editing performance and even model collapse.
@@
    We analyze the factors that contribute to superimposed noise. This understanding enables us to identify key mechanisms for reducing superimposed noise and enhancing the performance of sequential editing methods.
@@
    We develop DeltaEdit, an innovative sequential editing method that incorporates a dynamic orthogonal constraint strategy. Through extensive experiments, we demonstrate that DeltaEdit significantly outperforms existing methods.
Figure 1: Output deviation problem of edited LLM. The correctly activated output (green part) is interfered with by the incorrectly activated output (red parts).
@@ 2.1 Knowledge Editing
From the perspective of whether to modify model parameters, yao2023editinglargelanguagemodels categorizes knowledge editing methods into two major types: parameter-preserving methods and parameter-modifying methods. This paper primarily focuses on the latter. One line of work employs meta-learning approaches to edit language models through a hypernetwork. KE (decao2021editingfactualknowledgelanguage) utilizes a bidirectional LSTM to predict weight updates for editing, whereas MEND (mitchell2021fast) applies a low-rank decomposition of gradients to fine-tune the language model. Another line of work, based on conclusions drawn from causal probes (dai2022knowledgeneuronspretrainedtransformers; meng2023locatingeditingfactualassociations), performs edits in the feed-forward networks of middle model layers. KN (dai2022knowledgeneuronspretrainedtransformers) achieves knowledge editing by modifying the activation values of specific neurons. ROME (meng2023locatingeditingfactualassociations) uses normal equations to compute the update parameters required for editing, and MEMIT (meng2023masseditingmemorytransformer) further extends this approach to support batch editing. AnyEdit (jiang2025anyedit) effectively enables the editing of long-form knowledge.
@@ 2.2 Sequential Editing
gupta2024model points out that performing multiple consecutive edits can lead to model performance degradation. yang2024butterfly identified perplexity as an effective metric for detecting model collapse during sequence editing. By analyzing update parameters, hu2024wilkewiselayerknowledgeeditor finds that the overlap of input representations in the whitening space is a key factor contributing to poor editing performance. ma2024perturbationrestrainedsequentialmodelediting theoretically analyzes that the bottleneck limiting sequential editing in models lies in the condition number of matrices and proposes PRUNE, which supports sequential editing by controlling the growth of the condition number. gu2024modeleditingharmsgeneral observes that editing causes excessive parameter changes and proposes RECT, which improves editing performance by sparsifying update parameters. cai2024edit proposed O-Edit, which reduces interference between successive updates by orthogonalizing the direction of each knowledge update. fang2024alphaedit proposes AlphaEdit, which achieves nearly lossless sequential editing by restricting the solution space of update parameters to a specific null space.
3 Analysis on Sequential Editing
@@ 3.1 Preliminary
Autoregressive Language Model. Autoregressive language models (minaee2024large) generate the next token based on the preceding tokens. Modern autoregressive models are predominantly composed of multiple stacked transformer layers. In such models, the hidden state of the $l$ layer, denoted as $h^{l}$, is calculated as:
@@
Here, $a_{l}$ is the output of the attention block, $m_{l}$ is the output of the feed-forward network (FFN), $\sigma$ is the activation function, and $\gamma$ denotes layer normalization. Studies (geva2021transformerfeedforwardlayerskeyvalue) show that the parameter matrices of the FFN, $W_{\text{out}}$ and $W_{\text{in}}$, encode knowledge acquired during pretraining. Building on this finding, most knowledge editing methods focus on modifying the FFN to update knowledge.
Sequential Editing. Sequential editing (wang2024wiserethinkingknowledgememory) refers to performing continuous model editing on a model. Model editing aims to modify the knowledge encoded in a model. Following previous research (meng2023locatingeditingfactualassociations), we define knowledge in the form of triples, represented as $(s,r,o)$, where $s$ is the subject, $r$ is the relation, and $o$ is the object. For example, if a model memorizes the knowledge triple $(s=\text{iPhone},r=\text{Latest model},o=\text{iPhone 16})$, providing a language prompt $p(s,r)$, such as “The latest model of iPhone is,” as input causes the model to output “iPhone 16.” Model editing refers to the process of replacing the original knowledge triple $(s,r,o)$ with a new knowledge triple $(s,r,o^{*})$, and we denote this operation as $\mathcal{E}=(s,r,o,o^{*})$. In a sequential editing task of length $T$, where an editing sequence $\mathbb{E}_{T}=(\mathcal{E}_{1},\mathcal{E}_{2},\dots,\mathcal{E}_{T})$ is performed, each edit $\mathcal{E}_{i}$ builds upon the results of the previous edit $\mathcal{E}_{i-1}$.
Current knowledge editing methods (meng2023locatingeditingfactualassociations; meng2023masseditingmemorytransformer; fang2024alphaedit) primarily adopt the locate-then-edit paradigm to implement model editing. For a model $M$, these methods first locate the most appropriate parameters $W$ within the model, and then update them using the following approach:
@@ 3.2 Update Parameter
Most knowledge editing methods solve for update parameter $\Delta$ by applying the normal equation. For example, MEMIT (meng2023masseditingmemorytransformer) leverages the knowledge storage properties of FFN layers, identifies the $W_{out}^{l}$ of a suitable layer $l$ as the target for editing, and proposes the following equation:
@@
AlphaEdit (fang2024alphaedit) has a similar solution:
@@
$\Delta=\alpha\beta^{\top},\alpha=R,$ (9)
@@
Figure 2: The changes in Efficacytop and $noise_{E}$ with the number of edits. The left figure displays results for GPT2-XL using AlphaEdit and MEMIT, the right figure shows results for LLaMA3-8B using only AlphaEdit. MEMIT is excluded from the figure of LLaMA3-8B due to its excessively high $noise_{E}$, making visualization difficult.
@@ 3.3 The Problem of Superimposed Noise
Superimposed Noise. In the sequential editing task, a series of editing operations is performed on the model. Upon completing the entire sequence of edits $\mathbb{E}_{T}$, the parameters $W_{out}^{l}$ of the model’s $l$-th layer are updated. For the representation $k_{e}$ of the subject token being edited during operation $\mathcal{E}_{e}$, the output deviates from the original $W_{out}^{l}k_{e}$. This deviation can be formally analyzed using the $L_{2}$ norm:
@@
Here, $\|W_{out}^{l}k_{e}\|_{2}$ is a constant, while $\|\sum_{i}\Delta_{i}k_{e}\|_{2}$ determines the upper bound of the deviation. In editing operation $\mathcal{E}_{e}$, the update parameter can be expressed as $\Delta_{e}$. Ideally, after multiple sequential edits, different editing operations should remain independent and not interfere with one another. Under such conditions, the deviation associated with $\mathcal{E}_{e}$ satisfies:
@@
However, in practice, interference between different editing operations often occurs, leading to additional deviation in the output. To quantitatively characterize this phenomenon, we introduce the following definition.
Definition 1. The superimposed noise experienced by the editing operation $\mathcal{E}_{e}$, denoted as $noise_{e}$, is defined as
@@
This definition quantitatively describes the additional deviation, termed superimposed noise, which is induced by the interference among multiple editing operations. It provides a rigorous metric for evaluating the extent of such interference in sequential editing scenarios. To measure the overall superimposed noise level of all editing operations, we further define the average $noise_{E}$ of all editing operations as:
@@
The Impact of Noise on Editing Performance  To investigate the impact of superimposed noise on editing performance, experiments are conducted using AlphaEdit and MEMIT on GPT2-XL (Radford2019LanguageMA) and Llama3-8B (dubey2024llama). In these experiments, the models are sequentially edited using the CounterFact dataset (meng2023locatingeditingfactualassociations). Details of the experimental setup are provided in the Appendix A. The results presented in Figure 2 reveal the following trends:
@@
Moreover, as demonstrated in the Appendix C.6, models edited by MEMIT with high levels of superimposed noise lose the ability to generate coherent outputs. Unchecked accumulation of noise not only diminishes the efficacy of editing operations but can also substantially degrade overall model performance. Therefore, mitigating superimposed noise is critical for ensuring robust and reliable model behavior.
@@ 3.4 Factors Affecting Superimposed Noise
The experimental results are shown in Figure 3. The results indicate that the $k^{\top}\beta$ values obtained using AlphaEdit are significantly smaller than those of MEMIT. The improvement is primarily attributed to AlphaEdit’s use of null space projection and the inclusion of $K_{p}$ in the computation of $\beta$, which reduces the overlap of information between $\beta$ and the input representations of unrelated tokens. This demonstrates that optimizing $\beta$ to minimize erroneous activations is an effective strategy. However, as the number of sequential edits increases, the performance of AlphaEdit still degrades significantly. The observation suggests that simply reducing $k_{e}^{\top}\beta_{i}$ ($i\neq e$) is not sufficient to completely resolve the noise issue in sequential editing.
Figure 3: The variation curve of $k^{\top}\beta$ for AlphaEdit and MEMIT as the number of edits increases, with GPT2-XL serving as the edited model. As the results of Llama3-8B are not suitable for display in a line chart, they are presented in Appendix C.2
4 DeltaEdit
The analysis in Section 3.4 reveals that to further reduce superimposed noise, greater attention should be given to the other influencing factor $\alpha_{i}^{\top}\alpha_{j}$. To ensure that $\alpha_{i}^{\top}\alpha_{j}$ (for all $i,j,i\neq j$) approaches zero, it is essential to minimize the overlap of $\alpha$ with previous edits during training. Given the constrained solution space for $\alpha$, we propose DeltaEdit, a method that leverages a dynamic orthogonal constraints strategy based on historical editing information to reduce interference while ensuring editing efficiency.